OpenStudy (anonymous):

What is the sum of the coefficients of ( [3x - 3x^2 +1]^744 ) x ( [- 3x + 3x^2 +1]^745 ) ??

OpenStudy (anonymous):

i bet we can do this, i need paper

OpenStudy (anonymous):

Good problem.

OpenStudy (anonymous):

Thanks I Know Its Hard

OpenStudy (anonymous):

hmm i was thinking along the lines of \[(1+a)(1-a)=1-a^2\] with \[a=x^2-3x\] but that did't get me too far yet

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oh hold the phone. sum of coefficients is what you get when x = 1

OpenStudy (anonymous):

wow your smart :)

OpenStudy (anonymous):

so we can write as \[ \[1-9 (x^2-x)^2\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

that is \[(3x - 3x^2 +1 ) \times(- 3x + 3x^2 +1)=1-9(x^2-x)^2\] right?

OpenStudy (anonymous):

then raise to the power of whatever, say 755 and get \[(1-9(x^2-x)^2)^{755}\]

OpenStudy (anonymous):

how am i doing so far?

OpenStudy (anonymous):

wow :O your smart thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

replace x by 1, get 1

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (anonymous):

oh wait you have another term but you can cope with that

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[(1-9(x^2-x)^2)^{744}(- 3x + 3x^2 +1)\] is what you have

OpenStudy (anonymous):

replace x by 1, first term gives 1, second one gives ... 1 as well so i guess the answer is 1

OpenStudy (anonymous):

probably was an easier way to do it. repost and i bet you will get one

OpenStudy (anonymous):

thanks :)